OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 12x + 3y > 15. slope = 4, y-intercept = 5 slope = −4, y-intercept = 15 slope = -1/4, y-intercept = 5 slope = −4, y-intercept = 5

OpenStudy (anonymous):

You wan tot know how to solve it or do you just want the answer?

OpenStudy (anonymous):

both so i could learn next time

OpenStudy (anonymous):

To solve these aren't to hard right now you have 12x+3y>15, first the formula for the slope and y intercept equation is y=mx+b so we just have to move things around to get this. so subtract 12x from both sides and we get 3y=15-12x then divide both sides by three to get y=5-4x and your answer is y=-4x+5 in this equation y=mb+b, m is your slop and b is your why intercept Make sense?

OpenStudy (anonymous):

slope is -4 and y intercept is 5? if yes then thannk you!

OpenStudy (anonymous):

indeed did you get the process?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

Great